# Ageing-aware Energy Management for Residential Multi-Carrier Energy Systems

### v3

@@ 5.1 Case Study I: Day-ahead planning
The simulation workflow is presented in Algorithm 2 and depicted in Fig. 2. First, the hyperparameters are initialized. This includes the time window to be optimized $\mathcal{D}_{t}^{\pi}=[t;t+H]$, the number of days $n_{d}$, user preferences, the initial state $S_{a,0}$, and weights $w$. In our case for day-ahead planning, the time window is $H=48\text{hr}-\Delta t$ and $n_{d}=29$. The weights are $w_{\text{grid}}=w_{\text{loss}}=1$ and $w_{\text{SoCDep}}=w_{\text{TESS}}=1000$, since the first two represent real economic costs and the second represent penalties. At timestep $t$ , the OCP in Eq. 8, is solved, obtaining the action sequence $\mathcal{P}_{a,[t,t+24\text{hs}]}^{*}$. Together with the exogenous information sequence $\mathcal{W}_{[t+1,t+24\text{hs}]}$ the actions are passed to the simulator $S_{a,t}^{M}$ to get the feedback state sequence $\mathcal{S}_{a,[t,t+24\text{hs}]}$. This feedback loop is repeated $n_{d}$ times.
@@
As a representative example, Fig. 6 presents the results for the proposed CPBDeg planner for a monthly period. It has the resulting power balances (electrical and thermal) and the use of the hybrid energy storage system (HESS). The electric ESS have daily cycles to minimize operating costs (energy arbitrage). This is particularly important for the EV since its mobility demand already establishes a daily periodicity. Thus, due to the EV’s battery pack size and its natural periodicity, it becomes the main electric storage of the system. This frees up the BESS for energy arbitrage, trying to capture price variations when possible within the power balance. Thus, price volatility incentivizes cycling. However, due to the SEI model, batteries are also pushed downwards to the minimum ageing state at $\underline{SoC}_{b,t}$. Hence, the dispatch contemplates a trade-off between the 2 parts of the objective $C_{\textrm{grid}}$ and $Q_{\textrm{loss}}$. On the thermal carrier, the $P_{\text{load},t}^{\text{th}}$ represents the heat losses of the building which have to be compensated by either ST, HP or TESS. In the summer, the losses during solar hours are almost zero and thus the ST charges the TESS. If the prices are negative enough then HP charges the TESS during $\lambda_{t}\leq 0$. In the winter, the TESS is charged with the HP early in the day when prices are low or during solar hours, in that order of priority, to later deliver heat to the demand at high price hours. In summary, during the summer price volatility is high with several hours with $\lambda_{t}\leq 0$, the EMS buys this energy to reduce costs. In winter, prices are less volatile and the load is higher, leading to fewer opportunities for arbitrage and overall higher costs.
The schedules of the HESS under the different planners are summarized in Figs. 7- 9. In the BESS, Fig. 7, the BNoDeg and CEmpDeg cycle the battery pack more often. This is because BNoDeg does not contemplate ageing, and the CEmpDeg equations do not relate min $Q_{\text{loss}}$ to $\underline{SoC}$. The CPBDeg planner cycles less frequently, concentrating the operation near $P=0$ and $\underline{SoC}$ to reduce the ageing of the BESS. This is true for both summer (top) and winter (bottom). Moreover, on many days the price variations are not large enough to afford ageing the BESS, thus CPBDeg chooses to maintain the $SoC$ below the benchmarks. The two highlighted days are days where the CPBDeg outperforms the benchmarks and the BESS prioritizes $C_{\text{grid}}$ over $Q_{\text{loss}}$. Finally the linear BNoDeg is the planner with the highest percentage of rejects, roughly 45% of the total time in the summer. This is because when the BNoDeg plans for high-power discharges its predicted $\tilde{SoC}_{b,t}$ deviates from the true $SoC_{b,t}$ of the simulator $S^{M}_{a,t}$, wrongly depleting the ESS early (lower bound $\underline{SoC}_{b,t}$).
@@
For the EV timeseries comparison, Fig.8, the user’s mobility requirement leads to similar timeseries for all the planners. In general, CPBDeg delays the charging to extend the V2G as much as possible, to avoid $\overline{SoC}_{\textrm{EV}}$. V2G mode is less frequent in the BNoDeg/CEmpDeg than in the CPBDeg because of this, and is valid in both standard seasons. This counterintuitive result is because the ageing PBROM show how $Q_{\text{loss}}$ is minimized by $\underline{SoC}$. Since $p_{\text{SoCDep}}$ requires high $SoC^{*}_{\text{dep}}$, this results in more V2G.
On the thermal side, presented in Fig. 9, the natural periodicity of the carrier is longer due to its thermal efficiencies and size. Hence, during the first week the initial $SoC_{\textrm{TESS},0}$ influences costs greatly. After the first week, the EMS has already steered the buffer to its desired setpoint. This means a high $SoC_{\textrm{TESS},t}$ during summer (high heat generation, low load) and a low setpoint during winter (low heat generation, high load). The high setpoint during summer entails 2 risks: overcharging the TESS (i.e., activating soft-constraint and curtailing ST) and not capturing negative prices prices due to past short-sightedness (TESS starting a day with a high $SoC$). Even though in the summer CPBDeg has the smallest ST curtailment the ageing models have a low overal impact. In other words, integrating ageing models does not change the overall TESS decisions.
@@
BNoDeg 16.01 77.33 206.9 249.2
CEmpDeg 15.22 77.45 216.4 236.7
CPBDeg 20.28 78.25 206.6 234.3
Table 1: Planner comparison cost summary.
The performance of each planner is summarized in Table 1. Starting with the grid cost $C_{\textrm{grid}}$, the best performer is the CEmpDeg in the summer and BNoDeg in the winter. The differences between the first and second are less than 1 €. The worst performer is CPBDeg. For the total capacity fade $Q_{\textrm{loss}}$, CEmpDeg has the highest degradation in the summer and BNoDeg has the highest in the winter. The proposed ageing-aware CPBDeg achieves the lowest degradation in both seasons, maintaining a reasonable $C_{\textrm{grid}}$.
In general, the planner that has the lowest $C_{\textrm{grid}}$ is the one with the highest $Q_{\textrm{loss}}$, but this is not always BNoDeg. The CEmpDeg fails to minimize the total capacity fade because of its model bias ($S^{M}_{a,t}\neq\tilde{S}^{M}_{a,t}$), in which it is rewarded to cycle at $\overline{SoC}_{b}$ and a calendar ageing independent of the $SoC_{b,t}$. Additionally, the linear BNoDeg planner fails to fulfill its predictions because of the high number of rejected actions during summer.
@@
To dive deeper into the total storage usage, a quantitative analysis of the number of cycles done by the $b$ is necessary. Figure 10 presents the full equivalent cycles $FEC$ over time $t$, showing that the CPBDeg planners increase the EV usage $FEC_{\textrm{EV}}$ at the expense of doing fewer cycles with the BESS. When analyzing the capacity fade $Q_{\textrm{loss}}$ against the full eq. cycles $FEC$, Fig. 10, it is clear that the relative degradation per cycle ($\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$) of the CPBDeg is smaller in the winter than the used benchmarks. In summer, the BNoDeg has the lowest degradation per cycle due to its high number of rejects. The use of the EV s is the same for all planners, because the trajectory is mostly driven by user mobility. Finally, in both seasons, CEmpDeg presents the highest $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$. This appears to be a risky strategy due to a lack of consistency across seasons and objectives (minimizing degradation or minimizing grid costs). Lastly, even though the capacity fade is not significant in $T=1$ month, daily optimization can have a significant impact in the long term, as it was shown in [47, 46, 39]. As a final note, if the C-rate is increased ($\geq$1) and battery temperature $T$ is not constant, the degradation on a daily basis can be significantly higher.
@@
Finally, the computational time for the different strategies is presented in Fig. 11. Each sample is the total computational time it takes to solve Algorithm 2. Unexpectedly, BNoDeg has the lowest and most consistent $t_{\textrm{comp}}$ distribution, i.e., the smallest standard deviation. However, both CEmpDeg and CPBDeg planners have similar empirical distributions, maintaining overall fast computational time. Hence, the increase in modeling accuracy of PBROM is not prohibitively expensive when compared to its empirical counterpart. This is to be expected as the empirical ageing model is also non-linear and non-convex. In summary, this case study shows that:
@@
    CPBDeg achieves the smallest $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ (best ageing control) and smallest total $Q_{\text{loss}}$.
@@
    The computational cost of CPBDeg is not prohibitive when compared to its empirical benchmark CEmpDeg.
@@
    BNoDeg achieves one of the smallest $C_{\text{grid}}$ with fast computational times, as expected.
5.2 Case Study II: Manageing different cathodes
To demonstrate the PB models’ flexibility and extended capabilities, the CPBDeg scheduler is tested using two similar battery packs of the same rated capacity $Q$ but using different cells. One is formed with NMC cells and the other with LFP. Since LFP cells have a lower rated capacity of $Q_{n}=2.3$ Ah and a lower $OCV$, the battery packs have more $N_{s/p}$ to have roughly the same pack-rated capacity as their NMC counterparts. The power limits $\overline{P}_{b,t},\ \underline{P}_{b,t}$ are also maintained to make an even comparison.
@@ 5.2 Case Study II: Manageing different cathodes
The PB ageing models are suitable for both because they have graphite anodes [30, 15]. Nevertheless, they have different electrolytes. This is addressed by changing the electrolyte parameters in the model. The same model equations are used, but with different parameter values. This is a great advantage compared to the empirical fits presented in the literature. In the latter, the derived models are prone to overfitting to training conditions, delivering complex non-linear equations that can only be applied to specific chemistries and operating conditions.
@@
The simulation results are presented in Figs. 12 - 14. Starting with the EV, Fig. 12, the operation is similar except for a few days in summer and winter in which the LFP decides to have deeper discharges than its NMC counterpart. Moving forward to the BESS, Fig. 13, in summer, the LFP scheduler is too ambitious, generating early storage depletion roughly 40% of the time, 10% more than the NMC cells. During the winter, total rejections are reduced to 25% and 10% for LFP and NMC, respectively.
@@
Continuing with the ageing analysis Fig. 14 presents the $FEC$ and $Q_{\textrm{loss}}$ results. When looking at the EV, the $FEC_{\textrm{EV}}$ increases with the LFP cells. The opposite happens to the BESS, which reduces its cycles in the LFP pack. In combination with its lower calendar ageing, represented in the parameter set, the LFP packs achieve lower degradation per eq. cycle $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ than their NMC counterpart. As such, the CPBDeg-LFP reduces BESS total throughput ($FEC$) and capacity fade $Q_{\textrm{loss}}$ while increasing EV throughput and reducing its degradation.
Table 2 presents the summary of performance for both cell types. Overall, the CPBDeg-LFP achieves lower total costs and ageing despite its high number of rejected actions. During winter $C_{\textrm{grid}}$ stays the same with 37% less degradation. However, in the summer, just by changing the cell type from NMC to LFP $C_{\textrm{grid}}$ is reduced by 25%, similar to BNoDeg in Section 5.1, and roughly 25% less capacity fade. This shows that CPBDeg rightly exploits its physical information of the system to achieve better performance.
$C_{\text{g}}$ [€] $Q_{\text{loss}}$ [mAh]
@@
LFP 15.25 78.36 152.1 147.1
NMC 20.28 78.25 206.6 234.3
Table 2: Cathode comparison cost summary.
5.3 Case Study III: Manageing aged and fresh batteries

### v4

@@ 5.1 Case Study I: Degradation model comparison
Battery Scheduler Simulator
model BNoDeg CEmpDeg CPBDeg PBROM
Performance Bucket model ECM ECM SPM
Eq. 21 Eq. 21-23 Eq. 21-23 PBROM [Planden2022]
Ageing - Empirical PBROM PBROM
Eq. 24-25 [Wang2014] Eq. 25-29 [Jin2022] Eq. 25-29 [Jin2022]
Feature
Dynamic Only ECM ✓
Profile performance. Emp. Deg ✗
Cathode No distinction Only NMC
Chemistry
Used and Used cells are
New cells only smaller cells.
Table 2: Summary of battery models in each planner.
@@
A summary of the key Eqs. and features of each of the planners can be found in Tabel 2. The simulation workflow is presented in Algorithm 2 and depicted in Fig. 2. First, the hyperparameters are initialized. This includes the time window to be optimized $\mathcal{D}_{t}^{\pi}=[t;t+H]$, the number of days $n_{d}$, user preferences, the initial state $S_{a,0}$, and weights $w$. In our case $n_{d}=29$, the weights are $w_{\text{grid}}=1,\ w_{\text{loss}}=0.01$ and $w_{\text{SoCDep}}=w_{\text{TESS}}=1000$, since the first two represent real economic costs and the second represent penalties. At timestep $t$ , the OCP in Eq. 8, is solved, obtaining the action $P_{a,t}^{*}$. Together with the exogenous information $W_{t+1}$ the actions are passed to the simulator $S_{a,t}^{M}$ to get the feedback state $S_{a,t}$.
Figure 7: Monthly dispatch of MCES under CPBDeg controller with $w_{\text{loss}}=0.01$ for summer (left) and winter (right).
@@
As a representative example, Fig. 7 presents the results for the proposed CPBDeg controller for a monthly period. It has the resulting power balances (electrical and thermal) and the use of the hybrid energy storage system (HESS). The electric ESS have daily cycles to minimize operating costs (energy arbitrage). This is particularly important for the EV since its mobility demand already establishes a daily periodicity. Thus, due to the EV’s battery pack size and its natural periodicity, it becomes the main electric storage of the system. This frees up the BESS for energy arbitrage, trying to capture price variations when possible within the power balance. Thus, price volatility incentivizes cycling. However, due to the SEI model, batteries are also pushed downwards to the minimum ageing state at $\underline{SoC}_{b,t}$. Hence, the dispatch contemplates a trade-off between the 2 parts of the objective $C_{\textrm{grid}}$ and $Q_{\textrm{loss}}$. On the thermal carrier, the $P_{\text{load},t}^{\text{th}}$ represents the heat losses of the building which have to be compensated by either ST, HP or TESS. In the summer, the losses during solar hours are almost zero and thus the ST charges the TESS. If the prices are negative enough then HP charges the TESS during $\lambda_{t}\leq 0$. In the winter, the TESS is charged with the HP early in the day when prices are low or during solar hours, in that order of priority, to later deliver heat to the demand at high price hours. In summary, during the summer price volatility is high with several hours with $\lambda_{t}\leq 0$, the EMS buys this energy to reduce costs. In winter, prices are less volatile and the load is higher, leading to fewer opportunities for arbitrage and overall higher costs.
@@
The schedules of the HESS under the different controllers are summarized in Figs. 8- 10. In the BESS, Fig. 8, the BNoDeg and CPBDeg cycle the battery pack more often. This is because BNoDeg does not contemplate ageing, and the CPBDeg equations relate minimizing $Q_{\text{loss}}$ to $\underline{SoC}$. The CEmpDeg controller cycles less frequently, due to its overestimation of $Q_{\text{loss}}$, concentrating the operation around $\overline{SoC}$ to reduce the ageing of the BESS. This happens both in summer (top) and winter (bottom). Moreover, on many days the price variations are not large enough to afford ageing the BESS, thus CPBDeg chooses to maintain the $SoC$ close to its lower bound. The two highlighted days are days where the CPBDeg outperforms the benchmarks and the BESS prioritizes $C_{\text{grid}}$ over $Q_{\text{loss}}$.
For the EV timeseries comparison, Fig.9, the user’s mobility requirement leads to similar timeseries for all the planners. As expected, BNoDeg prioritizes short-term revenues and has the most frequent V2G, followed by CPBDeg and with CEmpDeg usually choosing the highest $SoC_{\text{EV},t}$ of the three. Again, this reflects the dependency of the empirical model, Eq. 23, with $DoD$. V2G mode suits the CPBDeg because it incentivizes discharging the EV, minimizing its degradation. This counterintuitive result is because the ageing PBROM shows how $Q_{\text{loss}}$ is minimized by $\underline{SoC}$. Since $p_{\text{SoCDep}}$ requires high $SoC^{*}_{\text{dep}}$, the dynamic combination of them both results in more V2G.
@@
On the thermal side, presented in Fig. 10, the natural periodicity of the carrier is longer due to its thermal efficiencies, size and C-rate. Qualitatively, all three planners take similar decisions independent of the battery model used. This is because the TESS has a roundtrip efficiency $\eta$ lower than the EV and BESS, and thus a longer optimization window $H$ is needed for the TESS to show its flexibility value relative to the BESS or EV [Darivianakis2017AManagement, Prat2024HowProblems]. During the first week the initial $SoC_{\textrm{TESS},0}$ influences costs greatly. After the first week, the EMS has already steered the buffer to its desired setpoint. This means a high $SoC_{\textrm{TESS},t}$ during summer (high heat generation, low load) and a low setpoint during winter (low heat generation, high load). The high setpoint during summer entails 2 risks: overcharging the TESS (i.e., activating soft-constraint and curtailing ST) and not capturing negative prices prices due to past short-sightedness (TESS starting a day with a high $SoC$). Even though in the summer CPBDeg has the smallest ST curtailment the ageing models have a low overal impact. In other words, integrating ageing models does not change the overall TESS decisions, since the advanced cell models don’t change the fact that $\eta_{b}\geq\eta_{\text{TESS}}$ and a bigger $H$ is needed for the TESS to present valuable flexibility.
@@
BNoDeg 55.4 101.7 377.3 400.1
CEmpDeg 62.3 115.3 234.0 250.8
CPBDeg 57.6 109.9 343.2 351.7
Table 3: Planner comparison cost summary with weights $w_{\text{grid}}=1,\ w_{\text{loss}}=0.01,\ w_{\text{SoC}}=w_{\text{TESS}}=1000$.
The performance of each planner is summarized in Table 3. Starting with the grid cost $C_{\textrm{grid}}$, the best performer is the BNoDeg in the summer and winter. The worst performer is CEmpDeg and the PB controller CPBDeg stands in the middle. For the total capacity fade $Q_{\textrm{loss}}$, CEmpDeg has the lowest degradation while the BNoDeg has the highest. The proposed ageing-aware CPBDeg has -10% less degradation than the BNoDeg, maintaining a reasonable $C_{\textrm{grid}}$.
@@
However, just reviewing the objective function $J^{\text{DA/MPC}}$ is not enough. To dive deeper into the total storage usage, a quantitative analysis of the number of cycles done by the $b$ is necessary. Figure 11 presents the full equivalent cycles $FEC$ over time $t$, showing that the BNoDeg controller has the most $FEC$, followed by CPBDeg and with the empirical controller CEmpDeg having the least throughput. When analyzing the capacity fade $Q_{\textrm{loss}}$ against the full eq. cycles $FEC$, Fig. 11, it is clear that the relative degradation per cycle ($\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$) of the CPBDeg is the smallest of them all. Moreover, its empirical competitor (CEmpDeg) has the highest degradation per cycle $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$, having the most inefficient degradation control. This appears to be a risky strategy due to a lack of consistency across seasons and objectives (minimizing degradation or minimizing grid costs). Lastly, even though the capacity fade is not significant in $T=1$ month, daily optimization can have a significant impact in the long term, as it was shown in [Movahedi2024ExtraMechanism, Cao2020MultiscaleModels, Reniers2021]. As a final note, if the C-rate is increased ($\geq$1) and battery temperature $T_{b,t}$ is not constant, the degradation on a daily basis can be significantly higher.
@@
BNoDeg CEmpDeg CPBDeg
# Variables 4416 6720 6720
# Constraints 3073 5377 5377
# Linear Eq. 2113 4033 4033
# Quad. Eq. 768 768 768
# Nonlin. Eq. 0 384 384
# Linear Ineq. 192 192 192
Table 4: Problem size and variables after discretization for each planner.
Finally, the computational time for the different strategies is presented in Fig. 12. Each sample is the total computational time it takes to solve Algorithm 2. All OCP instances are solved until local optimality, with BNoDeg being the closest to its global optima since its a convex QCQP [KNITRO]. Unexpectedly, BNoDeg has the lowest and most consistent $t_{\textrm{comp}}$ distribution, i.e., the smallest standard deviation. Both CEmpDeg and CPBDeg planners have similar distributions, maintaining overall fast computational time between 1-5s. More importantly, all three distributions overlap with more than 50% of cumulative probability. The optimization model variables and equations after transcription are presented in terms of their number of variables in Table 4. Hence, the increase in modeling accuracy of PBROM is not prohibitively expensive when compared to its empirical counterpart. This is to be expected as the empirical ageing model is also non-linear and non-convex. All of these demonstrates the feasibility of implementing the CPBDeg in real-time applications.
In summary, this case study shows that:
@@
    CPBDeg achieves the smallest $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ (best ageing control) while maintaining close to best $C_{\text{grid}}$.
@@
    The computational cost of CPBDeg is not prohibitive when compared to its benchmarks.
@@
    BNoDeg achieves the smallest $C_{\text{grid}}$ with fast computational times, as expected.
5.2 Case Study II: Managing different cathodes
To demonstrate the PB models’ flexibility and extended capabilities, the CPBDeg controller is tested using two similar battery packs of the same rated capacity $Q$ but using different cells. One is formed with NMC cells and the other with LFP. Since LFP cells have a lower rated capacity of $Q_{n}=2.3$ Ah and a lower $OCV$, the battery packs have more $N_{s/p}$ to have roughly the same pack-rated capacity as their NMC counterparts. The power limits $\overline{P}_{b,t},\ \underline{P}_{b,t}$ are also maintained to make an even comparison. Another weight $w_{\text{loss}}$ is chosen to further highlight differences between cathode chemistries. The driving profile (availability and consumption) are different from the previous Section, thus small deviations in costs are expected.
@@ 5.2 Case Study II: Managing different cathodes
The PB ageing models are suitable for both because they have graphite anodes [Jin2017, Jin2022]. Nevertheless, they have different electrolytes. This is addressed by changing the electrolyte parameters in the model. The same model equations are used, but with different parameter values. This is a great advantage compared to the empirical fits presented in the literature. In the latter, the derived models are prone to overfitting to training conditions, delivering complex non-linear equations that can only be applied to specific chemistries and operating conditions (stationary loads, fixed states).
@@
The simulation results are presented in Figs. 13 - 15. Starting with the EV, Fig. 13, the operation is similar except for a few days in summer and winter in which the LFP decides to have deeper discharges than its NMC counterpart. In other words, the actions in the NMC case are more conservative since the NMC is more sensitive to the SEI degradation. Moreover, with the LFP cells the CPBDeg chooses to do V2G more often than with the NMC cell. These leads to the CPBDeg-LFP achieving lower grid costs $C_{\text{grid}}$ due to the increased throughput.
@@
Continuing with the ageing analysis Fig. 15 presents the $FEC$ and $Q_{\textrm{loss}}$ results. The overall the $FEC$ are increased with the LFP cells. In combination with its lower calendar ageing, represented in the parameter set, the LFP packs achieve lower degradation per eq. cycle $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ than their NMC counterpart. As such, the CPBDeg-LFP increases both BESS and EV total throughput ($FEC$) while reducing its degradation $Q_{\textrm{loss}}$. Overall the change in cathode and consequent electrolyte, reduces the SEI degradation rate resulting in a more effective degradation control.
Table 5 presents the summary of performance for both cell types. Overall, the CPBDeg-LFP achieves lower total grid costs and ageing. During winter $C_{\textrm{grid}}$ stays the same ($\approx$4%) while the LFP cells has 20% less degradation. However, in the summer, the LFP $C_{\textrm{grid}}$ is 10% smaller than its NMC counterpart, with $C_{\text{grid}}$ similar to BNoDeg in Section 5.1 Table 3. The LFP summer case also has also roughly 20% less capacity fade than its NMC benchmark. This shows that CPBDeg rightly exploits its physical information of the system to achieve better performance, both lower grid cost and lower capacity fade.
$w_{\text{loss}}=0.1$ $C_{\text{g}}$ [€] $Q_{\text{loss}}$ [mAh]
@@
LFP 57.9 112.0 186.7 170.5
NMC 64.7 116.9 243.0 210.2
Table 5: Cathode comparison cost summary with weights $w_{\text{grid}}=1,\ w_{\text{SoC}}=w_{\text{TESS}}=1000$.
5.3 Case Study III: Managing aged and fresh batteries